### v1

@@ 2.3.1 Persistence
We assume the reader is already familiar with homology, but we nevertheless give a crash course in persistence and refer the curious to [14, 21] for further details. Whereas classical homology is useful for studying topological features of data, it lacks the discernment to pick the correct scale to extract features from. Persistent homology remedies this shortcoming by studying nested sequences of sublevel sets called filtrations, thereby extracting features across many scales. Taking a manifold $\mathcal{M}$ and a nice function to filter it with, for example, the height function, yields the persistence module comprised of homology groups and linear maps induced by the inclusions between sublevel sets
$\ldots\to H(\mathcal{M}_{a_{i-1}})\to H(\mathcal{M}_{a_{i}})\to\ldots\to H(\mathcal{M}_{a_{j-1}})\to H(\mathcal{M}_{a_{j}})\to\ldots.$ (1)
Composing the maps between consecutive groups, we get a map between any two groups in the module. We say a homology class $\alpha\in H(\mathcal{M}_{a_{i}})$ is born at $\mathcal{M}_{a_{i}}$ if it is not in the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{i}})$. If $\alpha$ is born at $\mathcal{M}_{a_{i}}$, it dies entering $\mathcal{M}_{a_{j}}$ if the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{j-1}})$ does not contain the image of $\alpha$, but the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(M_{a_{j}})$ does. The persistence of $\alpha$ is the difference between the function values at its birth and its death. If the function is a Morse function on a manifold, then precisely one Betti number $\beta_{p}$ changes when the threshold passes a critical value. If the index of the corresponding critical point is $p$, then either a $p$-dimensional class is born, so $\beta_{p}$ increases by one, or a $(p-1)$-dimensional class dies, so $\beta_{p-1}$ decreases by one. We use a persistence diagram ${\rm Dgm}{({f})}$ to encode the birth-death information of all of the $p$-dimensional homology classes of $M$ arising from the sublevel set filtration induced by the filtering function $f$, where each birth/death pair becomes a point in $\mathbb{R}^{2}$. See Figure 4.
Figure 4: A torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$. In the figure open black circles depict points of $H_{0}$- persistence, the red round dots depict points in ordinary $H_{l}$-persistence, the arrows indicate cycles that live forever (black for $H_{0}$, red for $H_{l}$, blue for $H_{1}$, and green for $H_{l+1}$).
@@ 2.3.2 Extended persistence
There are some drawbacks to the topological summary given by standard persistence, most notably points at infinity. For example, consider a sublevel set filtration of a shape embedded in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$ with nontrivial $H_{0}$ and $H_{1}$: at some point, connected components and loops are born, but never die, as they are present in all sublevel sets after their initial appearance. Even worse, if the input is non-generic and two $H_{1}$ cycles are born at the same height in the sublevel set filtration, then they give rise to identical persistence pairs.
@@
To address this, Agarwal et. al. [1] established a pairing between all critical points of a height function on a 2-manifold, which Cohen-Steiner et. al. [10] extended to general manifolds with tame functions, leveraging Poincaré and Lefschetz duality to create a new sequence of homology groups where we begin and end with the trivial group. This guarantees that each homology class that is born will also die at a finite value, replacing all the problematic points paired with $\infty$, and guarantees a perfect matching on critical points. Let $H_{d}(\mathcal{M},\mathcal{M}^{a})$ denote the relative homology group of $\mathcal{M}$ with the superlevel set $\mathcal{M}^{a}=f^{-1}[a,\infty)$. Again assume we have a tame function and critical set $a_{1},\cdots,a_{k}$, and note that $H_{d}(\mathcal{M}_{a_{k}})=H_{d}(\mathcal{M})=H_{d}(\mathcal{M},\mathcal{M}^{a})$ for any $a>a_{k}$. From this, we can create a new sequence of homology groups
00 0 $\rightarrow H_{d}(\mathcal{M}_{a_{1}})\rightarrow H_{d}(\mathcal{M}_{a_{2}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M}_{a_{k-1}})$
$\rightarrow H_{d}(\mathcal{M}_{a_{k}})=H_{d}(\mathcal{M},\emptyset)\rightarrow\ H_{d}(\mathcal{M},\mathcal{M}^{a_{k}})$
$\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{k-1}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{1}})=H_{d}(\mathcal{M},\mathcal{M})=0,$ (2)
@@
which we call the $d^{th}$ extended filtration sequence. We define the sequence $H_{d}(\mathcal{M}_{a_{1}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M}_{a_{k}})$ as the upwards sequence and the sequence $H_{d}(\mathcal{M},\mathcal{M}^{a_{k}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{1}})$ as the downwards sequence.
@@
Note that this sequence fits the structure of a persistent module, so like standard persistence, it also has a unique interval decomposition. The only change is that we interpret the persistence points differently in this setting. Specifically, the points in the persistence diagram can be partitioned into three different groups: 1) the classes which are born and die in the upwards sequence, 2) the classes which are born and die in the downwards sequence, and 3) the classes which are born in the upwards sequence and die in the downwards sequence. Further, we associate the index of birth and death intervals with the value $a_{i}$ for both $H_{p}(\mathcal{M}_{a_{i}})$ in the the upwards sequence and $H_{p}(\mathcal{M},\mathcal{M}^{a_{i}})$ in the downwards sequence.
This creates three different classes of persistence pairs: those that correspond to a class that is born and dies in the upwards sweep (ordinary persistence points), those that correspond to a class that is born and then dies in the downwards sweep (relative persistence points), and those that correspond to a class that is born in the upwards sweep and dies in the downwards sweep (extended persistence points). We denote the class of ordinary, relative, and extended points as $\text{Ord}_{d}(f)$, $\text{Rel}_{d}(f)$, and $\text{Ext}_{d}(f)$, respectively. Note that the ordinary diagram is entirely above the diagonal; the relative diagram is entirely below the diagonal; while the extended persistence points can be on either side. See Figure 5.
Figure 5: A torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$. In the figure open black circles depict points of ordinary or extended $H_{0}$- persistence, the red round dots depict points in ordinary $H_{l}$-persistence, the green squares (below the diagonal) indicate relative $H_{l+1}$-persistence and the blue crosses (below the diagonal) indicate relative $H_{1}$-persistence.
@@ 2.3.3 Vineyards
One can also study the persistence on a manifold $\mathcal{M}$ arising from multiple functions. As long as the functions are similar enough, the Stability Theorem [9] of persistent homology asserts that their associated persistence diagrams will also be similar; see also [28] for a discussion of the algebraic details of maps between “nearby" diagrams. More precisely, the bottleneck distance between two persistence diagrams of functions $g_{u},g_{v}\colon M\to{\mathbb{R}}$ is bounded from above by the infinity norm of their difference:
${W_{\infty}}(\mathscr{D}_{\cdot}(g_{u}),\mathscr{D}_{\cdot}(g_{v}))$ $\leq\|{g_{u}}-{g_{v}}\|_{\infty},$ (3)
@@
in which $\mathscr{D}_{\cdot}(g)$ is the persistence diagram of $g$, a matching between the two diagrams is quantified by the supremum of the max-distances between matched points, and ${W_{\infty}}$ is the infimum over all possible matchings, in which we allow the introduction of points on the diagonal (where birth time is equal to death time), which we preferably match with points near to the diagnal to points on the diagonal. Intuitively, the bottleneck distance describes the worst disparity between the best matching of points in persistence diagrams: the worst disparity is the “bottleneck” preventing a smaller distance.
This gives rise to the concept of a vineyard [11, 28]. It formalizes the idea that a feature of $g_{u}$ is still recognizable in $g_{v}$, provided the two functions $u$ and $v$ are not too far apart. Features are points in the diagram, and the association is a matching between the points of $g_{u}$ and of $g_{v}$. Cohen-Steiner et al. [11] propose the so-called vineyard algorithm that traces the features (points) while continuously deforming $g_{u}$ into $g_{v}$, realized as an update to the reduced matrix arising from $g_{u}$. This update can be done in $O(n)$ time, as compared to the worst case $O(n^{3})$ time to compute and reduce the new matrix from scratch.
3 Contribution
@@
In the context of persistence of the induced distance function, we need to introduce a number of conventions for monodromy to be well defined. In fact we will introduce monodromy both associated to an entire vineyard as well as to a single vine in the vineyard. We will then present our main results, demonstrating monodromy and braiding in detail through several examples.
3.1 Monodromy in vineyards: a geometric viewpoint
@@
Let $\mathcal{M}$ be a manifold embedded in $\mathbb{R}^{d}$ and $\gamma:[0,2\pi]\to\mathbb{R}^{d}$ be a parametrization of a loop $\gamma$. Let $d(\cdot,\gamma(t)):\mathcal{M}\to{\mathbb{R}}$ be the function $x\mapsto d(x,\gamma(t))_{\mathcal{M}}$, where $d(x,\gamma(t))$ is the Euclidean distance from $x$ to $\gamma(t)$ and, as indicated, $d(x,\gamma(t))_{\mathcal{M}}$ is its restriction to the manifold $\mathcal{M}$. We will now identify the ends of the interval $[0,2\pi]$ – that is, we pass to $\mathbb{S}^{1}$. For each $t$ the function $d(\cdot,\gamma(t))_{\mathcal{M}}$ induces a filtration on $\mathcal{M}$, by the sub-level sets of the function. Therefore we can consider (for each $t$) the $l$th order persistence diagram of this filtration $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$. The map
@@
$t$ $\mapsto(t,\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})),$
where $\mathrm{Dgm}$ is the space of persistence diagrams, is (trivially) a covering space of $\mathbb{S}^{1}$. We refer to $\operatorname{CV}_{\mathcal{M}}$ as the closed vineyard map. Thanks to [9] the points in the persistence diagram are (Lipschitz) continuous with respect to $t$. The map $\operatorname{CV}_{\mathcal{M}}$ is illustrated in Figure 6.
Figure 6: An illustration of the map $\operatorname{CV}_{\mathcal{M}}$. We indicate the persistence diagrams, which form the fibers, only explicitly in a number of places for visual clarity.
We will now define monodromy in two increasingly more complicated settings. Our definition is canonical only in the first setting, while the second definition involves a choice. We note that generically [28, page 3], there are no points of higher multiplicity in a persistence diagram and hence the connected components in the image of $\operatorname{CV}_{\mathcal{M}}$, that is the vines, are non-intersecting curves. We fix the assumption throughout this paper that our vineyards are generic in this sense.
Staying clear of the diagonal
This leads us to the definition of monodromy in the simplest setting (where we stay clear of the diagonal):
@@
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal.
@@
Individual vine Given a point in the persistence diagram $\mathscr{D}_{l}(d(\cdot,\gamma(\tau))_{\mathcal{M}})$ for some $\tau\in\mathbb{S}^{1}$, write $\mathcal{V}(t)$ for the lift of $\gamma(t)$, which continuously assigns a point in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$ for every $t$ and yields the given point for $t=\tau$. We will now assume without loss of generality that $\tau=0$. We say that the vine $\mathcal{V}(t)$ exhibits monodromy if the start and end point of the lifted curve do not coincide. Similar to $\mathcal{V}(t)$, write $\tilde{\mathcal{V}}^{k}(t)$ for the analogous lift of $\gamma^{k}(t)$. Then $\mathcal{V}(t)$ exhibits monodromy of order $k$ if $k$ is the smallest integer strictly larger than 00, such that $\tilde{\mathcal{V}}^{k}(0)=\tilde{\mathcal{V}}^{k}(2\pi k)$. In other words the order of the monodromy is the number of points above any $t$ the connected component in the image of $\operatorname{CV}_{\mathcal{M}}$ that contains the given point in $\mathscr{D}_{l}(d(\cdot,\gamma(\tau))_{\mathcal{M}})$.

### v2

@@ 2.3.1 Persistence
Topological data analysis (TDA) starts from the premise that while classical homology is useful for studying topological features of data, it lacks the discernment to pick the correct scale to extract features from. Persistent homology remedies this shortcoming by studying nested sequences of sublevel sets called filtrations, thereby extracting features across many scales. Taking a manifold $\mathcal{M}$ and a nice function to filter it with, for example, the height function, yields the persistence module comprised of homology groups and linear maps induced by the inclusions between sublevel sets
(2.1) $\ldots\to H(\mathcal{M}_{a_{i-1}})\to H(\mathcal{M}_{a_{i}})\to\ldots\to H(\mathcal{M}_{a_{j-1}})\to H(\mathcal{M}_{a_{j}})\to\ldots.$
Composing the maps between consecutive groups, we get a map between any two groups in the module. We say a homology class $\alpha\in H(\mathcal{M}_{a_{i}})$ is born at $\mathcal{M}_{a_{i}}$ if it is not in the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{i}})$. If $\alpha$ is born at $\mathcal{M}_{a_{i}}$, it dies entering $\mathcal{M}_{a_{j}}$ if the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{j}})$ does not contain the image of $\alpha$, but the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(M_{a_{j-1}})$ does. The persistence of $\alpha$ is the difference between the function values at its birth and its death. If the function is a Morse function on a manifold, then precisely one Betti number $\beta_{{\tilde{\ell}}}$ changes when the threshold passes a critical value. If the order of the corresponding critical point is $\ell$}, then either a $\ell$}-dimensional class is born, so $\beta_{\ell}$ increases by one, or a $({\ell}-1)$-dimensional class dies, so $\beta_{{\ell}-1}$ decreases by one. We use a persistence diagram ${\rm Dgm}{({f})}$ to encode the birth-death information of all of the $p$-dimensional homology classes of $M$ arising from the sublevel set filtration induced by the filtering function $f$, where each birth/death pair becomes a point in $\mathbb{R}^{2}$. See Figure 2.3.
Figure 2.3: We consider a torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$ with a height function (whose direction is indicated by the black arrow). In the figure, the color indicates the degree in homology of the point: black for $H_{0}$; light blue for $H_{1}$; red for $H_{l}$; and dark green for $H_{l+1}$. Arrow markers distinguish cycles that live forever from those with finite death coordinates.
@@ 2.3.2 Extended persistence
There are some drawbacks to the topological summary given by standard persistence, most notably from our perspective the points at infinity, which could cause difficulties in the next section when we formalize the notion of a braided vineyard. For example, consider a sublevel set filtration of a shape embedded in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$ with nontrivial $H_{0}$ and $H_{1}$: at some point, connected components and loops are born, but never die, as they are present in all sublevel sets after their initial appearance. Even worse, if the input is non-generic and two $H_{1}$ cycles are born at the same height in the sublevel set filtration, then they give rise to identical persistence pairs.
@@
To address this, Agarwal et. al. [1] established a pairing between all critical points of a height function on a 2-manifold, which Cohen-Steiner et. al. [14] extended to general manifolds with tame functions, leveraging Poincaré and Lefschetz duality to create a new sequence of homology groups where we begin and end with the trivial group. This guarantees that each homology class that is born will also die at a finite value, replacing all the problematic points paired with $\infty$, and guaranteeing a perfect matching on critical points. Let $H_{l}(\mathcal{M},\mathcal{M}^{a})$ denote the relative homology group of $\mathcal{M}$ with the superlevel set $\mathcal{M}^{a}=f^{-1}[a,\infty)$. Again assume we have a tame function and critical set $a_{1},\cdots,a_{k}$, and note that $H_{l}(\mathcal{M}_{a_{k}})=H_{l}(\mathcal{M})=H_{l}(\mathcal{M},\mathcal{M}^{a})$ for any $a>a_{k}$. From this, we can create a new sequence of homology groups
$0$ $\rightarrow H_{l}(\mathcal{M}_{a_{1}})\rightarrow H_{l}(\mathcal{M}_{a_{2}})\rightarrow\ldots\rightarrow H_{l}(\mathcal{M}_{a_{k-1}})$
$\rightarrow H_{l}(\mathcal{M}_{a_{k}})=H_{l}(\mathcal{M},\emptyset)\rightarrow\ H_{l}(\mathcal{M},\mathcal{M}^{a_{k}})$
(2.2) $\rightarrow H_{l}(\mathcal{M},\mathcal{M}^{a_{k-1}})\rightarrow\ldots\rightarrow H_{l}(\mathcal{M},\mathcal{M}^{a_{1}})=H_{l}(\mathcal{M},\mathcal{M})=0,$
@@
which we call the $l^{th}$ extended filtration sequence. We define the sequence $H_{l}(\mathcal{M}_{a_{1}})\rightarrow\ldots\rightarrow H_{l}(\mathcal{M}_{a_{k}})$ as the upwards sequence and the sequence $H_{l}(\mathcal{M},\mathcal{M}^{a_{k}})\rightarrow\ldots\rightarrow H_{l}(\mathcal{M},\mathcal{M}^{a_{1}})$ as the downwards sequence.
@@
Note that this sequence fits the structure of a persistence module, so like standard persistence, it has a unique interval decomposition [16]. The only change is that we interpret the persistence points differently in this setting. Specifically, the points in the persistence diagram can be partitioned into three different groups: 1) the classes which are born and die in the upwards sequence, which we denote as ordinary persistence points, 2) the classes which are born and die in the downwards sequence, which we call relative persistence points, and 3) the classes which are born in the upwards sequence and die in the downwards sequence, which we call the extended persistence points. Further, we associate the order of birth and death intervals with the value $a_{i}$ for both $H_{l}(\mathcal{M}_{a_{i}})$ in the the upwards sequence and $H_{l}(\mathcal{M},\mathcal{M}^{a_{i}})$ in the downwards sequence. We stress that we use the value $a_{i}$ as the (birth or death) coordinate in the persistence diagram for both $H_{l}(\mathcal{M}_{a_{i}})$ in the the upwards sequence and $H_{l}(\mathcal{M},\mathcal{M}^{a_{i}})$ in the downwards sequence. Note that this means points in the ordinary diagram are entirely above the diagonal; the relative points lie entirely below the diagonal; while the extended persistence points can be on either side. See Figure 2.4.
Figure 2.4: A torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$ (left) with a height function (whose direction is indicated by the black arrow). In the figure, the colour indicates the degree in homology of the point: black for $H_{0}$; light blue for $H_{1}$; red for $H_{l}$; and dark green for $H_{l+1}$; while the shape of the marker indicates the type of the point, dots for ordinary points, that is they are both born and die on the way up; squares for relative points, that is they are both born and die on the way down; and triangles for extended points, that is they are born upwards but die downwards.
@@ 2.3.3 Vineyards
One can also study the persistence on a manifold $\mathcal{M}$ arising from multiple functions. As long as the functions are similar enough, the Stability Theorem [13] of persistent homology asserts that their associated persistence diagrams will also be similar; see also [47] for a discussion of the algebraic details of maps between “nearby" diagrams. More precisely, the bottleneck distance ($W_{\infty}$}) between two persistence diagrams of functions $g_{u},g_{v}\colon\mathcal{M}\to{\mathbb{R}}$ is bounded from above by the infinity norm of their difference:
(2.3) ${W_{\infty}}({\rm Dgm}{({g_{u}})},{\rm Dgm}{({g_{v}})})$ $\leq\|{g_{u}}-{g_{v}}\|_{\infty},$
@@
in which ${\rm Dgm}{({g})}$ is the persistence diagram of $g$, a matching between the two diagrams is quantified by the supremum of the max-distances between matched points, and $W_{\infty}$} is the infimum distance over all possible matchings, where we allow the introduction of points on the diagonal (where birth time is equal to death time), in which we prefereably match points near to the diagonal to points on the diagonal. Intuitively, the bottleneck distance describes the worst disparity between the best matching of points in persistence diagrams: the worst disparity is the “bottleneck” preventing a smaller distance.
This gives rise to the concept of a vineyard [15, 47]. It formalizes the idea that a feature of $g_{u}$ is still recognizable in $g_{v}$, provided the two functions $u$ and $v$ are not too far apart. Features are points in the diagram, and the association is a matching between the points of $g_{u}$ and of $g_{v}$. We now give a brief overview of the definition, but refer the reader to [47] for a more detailed description. Given a one-parameter family of continuous (Morse) functions (or more generally filtrations) parametrized by $t\in I$ or $t\in\mathbb{S}^{1}$, with $I$ an interval in $\mathbb{R}$, the vineyard is a section of trivial fibre bundle (consisting of the product of the parameter domain ($I$ or $\mathbb{S}^{1}$) and the space of persistence diagrams), which associates to each parameter value the persistence diagram of the (sub/super) level set filtration induced by the Morse function. In our context (see below) there will be no points in persistence diagrams that have multiplicity higher than one, in which case the structure is simple. Each point on a persistence diagram extends to a Lipschitz curve in the vineyard thanks to the aforementioned stability of persistence diagrams. In other words each connected component of the section of the bundle is a Lipschitz curve. We call each such connected component a vine. We’ll come back to this definition in Section 3, when defining monodromy in our context and introduce additional notation.
We briefly note that vineyards have been used in several settings for analysis of dynamic data sets [52, 8, 50, 10]. Cohen-Steiner et al. [15] proposed the so-called vineyard algorithm that traces the features (points) while continuously deforming $g_{u}$ into $g_{v}$, realized as an update to the reduced matrix arising from $g_{u}$. This update can be done in $O(n)$ time, as compared to the alternative (in practice) $O(n^{3})$ time [^3] Formally speaking the complexity is that of matrix multiplication, i.e. $O(n^{\omega})$, for which the best currently known value is $\omega\simeq 2.371552$, see [49]. of computing and reducing the new matrix from scratch. The algorithm is available as a part of the package Dionysus [33].
3 Monodromy in vineyards: a geometric viewpoint
@@
In the context of persistence of some topological space with the induced distance function from a point, we need to introduce a number of conventions for monodromy to be well defined. In fact we will introduce monodromy both associated to an entire vineyard as well as to a single vine in the vineyard.
Definition 3.1 (Radial transform).
We follow the convention from [38] and define the radial transform as the map that associates to any point $p\in\mathbb{R}^{d}$ the diagram $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,p)|_{\mathcal{M}}))\in\mathrm{Dgm}$.
@@
We now consider a restriction of the radial transform. Let $\mathcal{M}$ be a manifold embedded in $\mathbb{R}^{d}$ and $\gamma:[0,2\pi]\to\mathbb{R}^{d}$ be a parametrization of a loop $\gamma$. Let $d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}}:\mathcal{M}\to{\mathbb{R}}$ be the function $x\mapsto d_{\mathbb{E}}(x,\gamma(t))$, where $d_{\mathbb{E}}(x,\gamma(t))$ is the Euclidean distance from $x$ to $\gamma(t)$ and, as indicated, $|_{\mathcal{M}}$ is used to emphasize that we exclusively consider $x\in\mathcal{M}$. We will now identify the ends of the interval $0,2\pi]$ – that is, we pass to $\mathbb{S}^{1}$. For each $t$ the function $d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}}$ induces a filtration on $\mathcal{M}$, by the sub-level sets of the function. Therefore we can consider (for each $t$) the $l$th order persistence diagram of this filtration $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$. The map
@@
$t$ $\mapsto(t,\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})),$
where $\mathrm{Dgm}$ is the space of persistence diagrams, is (trivially) a covering space of $\mathbb{S}^{1}$. We refer to $\operatorname{CV}_{\mathcal{M}}$ as the closed vineyard map. Thanks to [13] the points in the persistence diagram are (Lipschitz) continuous with respect to $t$. The map $\operatorname{CV}_{\mathcal{M}}$ is illustrated in Figure 3.1.
Figure 3.1: An illustration of the map $\operatorname{CV}_{\mathcal{M}}$. We indicate the persistence diagrams, which form the fibers, only explicitly in a number of places for reasons of visibility. The knot diagram below (black), define a ouroborus knot. Here, we define the ouroborus as an unknot whose knot diagram is a (finite or segment of a) spiral with the end points connected with an monotone curve (where monotonicity refers to the relation between the angular and radial coordinates) such that the knot diagram only has under or over crossings. The observation loop is indicated in blue.
We note that generically [47, page 3], there are no points of higher multiplicity in a persistence diagram and hence the connected components in the image of $\operatorname{CV}_{\mathcal{M}}$, that is the vines, are non-intersecting curves. We fix the assumption throughout this paper that our vineyards are generic in this sense.
This leads us to the definition of monodromy in the simplest setting (where vines stay clear of the diagonal), which nonetheless is all that is needed for our constructions. Intuitively, we want to codify monodromy in terms of an integer, which encodes how many copies of the vineyard must be glued together in order for points to return to their start. More formally:
@@
Assume that for all points $\gamma(t)$, there are no points of multiplicity higher than $1$ in $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$, which means that all the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal. We define the following:
@@
Individual vine monodromy Given a point in the persistence diagram $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(\tau))|_{\mathcal{M}})$ for some $\tau\in\mathbb{S}^{1}$, write $\mathcal{V}(t)$ for the lift of $\gamma(t)$, which continuously assigns a point in $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(t))|_{\mathcal{M}})$ for every $t$ and yields the given point for $t=\tau$. We will now assume without loss of generality that $\tau=0$. We say that the vine $\mathcal{V}(t)$ exhibits monodromy if the start and end points of the lifted curve do not coincide. Similar to $\mathcal{V}(t)$, write $\tilde{\mathcal{V}}^{k}(t)$ for the analogous lift of $\gamma^{k}(t)$. Then $\mathcal{V}(t)$ exhibits monodromy of order $k$ if $k$ is the smallest integer strictly larger than 0, such that $\tilde{\mathcal{V}}^{k}(0)=\tilde{\mathcal{V}}^{k}(2\pi k)$. In other words, the order of the monodromy is the number of points above any $t$ in the connected component in the image of $\operatorname{CV}_{\mathcal{M}}$ that contains the given point in $\mathscr{D}_{l}(d_{\mathbb{E}}(\cdot,\gamma(\tau))|_{\mathcal{M}})$.